# Some Aspects of Three-Quark Potentials (Part II)

### v1

@@ I Introduction
The paper is organized as follows. In Sec.II, we set the framework and recall some preliminary results. Then, in Sec.III we construct and analyze the three-quark potential for quarks located at the vertices of an isosceles triangle. We continue in Sec.IV with a collinear geometry, where the quarks lie on a straight line. Here we also discuss diquarks. In Sec.V, we consider the IR limit for a general quark configuration and explores the asymptotic behavior of the potential. We conclude in Sec.VI with a discussion of the implications of our findings and some open problems. Appendix A introduces our notation and lists some useful formulas. To make the paper more self-contained, we include technical details and derivations in Appendices B and C. Appendices D, E, and F provide supplementary material related to Sec.III-IV.
@@ II Preliminaries
Here $r$ denotes the fifth (radial) coordinate, and $\mathsf{s}$ is a deformation parameter. The boundary lies at $r=0$, while the soft wall at $r=1/\sqrt{\mathsf{s}}$. The wall effectively prevents strings from penetrating deep into the bulk.
@@ III Isosceles Triangle Geometry
which are valid under the restriction $20\leq\beta\leq(\pi-\theta)/$ 2.
@@ III.1.2 Configuration II
Now we would like to examine the $-\ell$ behavior of $E^{\,\text{\tiny II}}$. As before, the basic idea is to let one or both of the $\lambda$ parameters approach $1$. There are two essentially distinct cases to consider: $\lambda_{1}\rightarrow 1$ at fixed $\lambda_{3}<1$, where only strings 1 and 2 (ending on $Q_{1}$ and $Q_{2}$) become infinitely long, and $\lambda_{1},\,\lambda_{3}\rightarrow 1$ , where all the strings become infinitely long.
@@
The key new feature here is that the constant term depends explicitly on the apex angle $\theta$. This has a simple explanation. In the limit as $\lambda_{1}\rightarrow 1$, the two strings become infinitely long, while the third remains finite in size. Its length depends on the apex angle that directly affects the constant term but not the leading one.
@@ III.1.3 Configuration III
Finally, consider configuration $III$. A straightforward way to proceed is as follows. Since $\alpha_{3}$ changes its sign from negative to positive, the expressions for the geometrical constraint, leg length, and energy can be obtained from those of subsection 2 by respectively replacing ${\cal L}^{-}$ and ${\cal E}^{-}$ with ${\cal L}^{+}$ and ${\cal E}^{+}$, respectively. The force balance equations, however, remain unchanged. Thus, we can rewrite (3.8) and (3.9) as
@@ III.2.1 $\theta<\frac{2}{3}\pi$
The two main conclusions follow from this Table. First, as seen from the fourth column, the $Y$-point generally does not coincide with the Fermat point $F$ of the triangle. [^12] As shown in subsection A, $Y$ approaches $F$ in the limit $\ell\rightarrow\infty$. The only exception is the case $\theta=\frac{\pi}{3}$, where the two points coincide due to symmetry. Second, the last column shows that the coupling $\alpha_{\text{\tiny QQ}}$ depends only weakly on $\theta$, and that the ratio $\alpha_{\text{\tiny QQ}}/\alpha_{\text{\tiny Q}\bar{\text{\tiny Q}}}$ is very close to the phenomenologically expected value of $1/2$. Since the discrepancy among the $\alpha_{\text{\tiny QQ}}$ values does not exceed $1.5\%$, the assumption of pairwise quark interaction appears to be quite reasonable.
@@ IV.2 Some examples
where ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ is the solution to Eq.(4.8) (see Table II).
@@
where $C_{\text{\tiny 3Q}}=C^{\,\text{\tiny A}}$. Notably, the $\Delta$-law yields the same leading term. Since it also provides a good approximation at small $\ell$, one may expect that the difference between the genuine three-quark potential and the $\Delta$-law is small. To explore this, we compare the two in the right panel of the Figure, which plots the potential derived from the $\Delta$-law. The visible difference arises from the constant terms in the $-\ell$ expansions and amounts to approximately $57,\text{MeV}$, as indicated in Table II. This difference could potentially be determined with reasonable accuracy through numerical simulations.
@@ IV.3 A remark on diquarks
with $L^{\text{\,\tiny A,\tiny B}}=\frac{1}{\sqrt{\mathsf{s}}}\bigl{(}{\cal L}^{-}(\lambda_{2},v)-{\cal L}^{+}(\alpha_{3},v)\bigr{)}$. The bounds on $v$ are determined by the Equations: $L(v)=\ell$ for $\text{v}_{\ast}$, $\alpha_{1}(v)=1$ for ${\text{v}}_{\text{\tiny 0}}$, and $\lambda_{2}(v)=1$ for ${\text{v}}_{\text{\tiny 1}}$. The second corresponds to a transition between the configurations, while the third to the limit of infinitely long string ending on $Q_{2}$. In contrast to the constraint (4.2), the constraint (4.23) does not prevent us from taking the limit $\lambda_{2}\rightarrow 1$ ($L\rightarrow\infty$) for configuration B.
@@ Appendix C Some details on the quark-antiquark potential
Figure 14: A string configuration in five dimensions. Here, $\ell$ denotes the distance between the two quarks, and $r_{\text{\tiny 0}}$ is the radial coordinate of the turning point at $x=0$.
@@ D.3 The range $\theta^{1}\leq\theta\leq\theta^{3}$
whose zeros correspond to the solutions of $\alpha_{3}(v)=0$ if $\beta(v)$ and $\alpha_{1}(v)$ are given by the first two equations in (3.33). As before, we are interested in the zeros within the interval $0<v<{\text{v}}_{\text{\tiny 1}}$. A numerical analysis of $f^{3}$ is straightforward and shows two distinct cases: the function has no zeros for $\theta^{1}\leq\theta<0.350\pi$, and has two zeros for $0.350\pi\leq\theta\leq\theta^{3}$. The former case aligns with the earlier discussion of $\theta=\frac{\pi}{3}$, while the latter implies transitions between configurations II and III. For illustration, we plot $f^{3}$ as a function of $v$ in Figure 17. The potential is described
@@
in terms of these configurations as follows:

### v2

@@ I Introduction
The paper is organized as follows. In Sec.II, we set the framework and recall some preliminary results. Then, in Sec.III we construct and analyze the three-quark potential for quarks located at the vertices of an isosceles triangle. We continue in Sec.IV with a collinear geometry, where the quarks lie on a straight line. Here we also discuss diquarks. In Sec.V, we consider the IR limit for a general quark configuration and explores the asymptotic behavior of the potential. We conclude in Sec.VI with a discussion of the implications of our findings and some open problems. Appendix A introduces our notation and lists some useful formulas. To make the paper more self-contained, we include technical details and derivations in Appendices B and C. Appendices D, E, and F provide supplementary material related to Sec.III-V.
@@ II Preliminaries
Here $r$ denotes the fifth (radial) coordinate, $\mathsf{s}$ is a deformation parameter, and $i=1,2,3$. The boundary lies at $r=0$, while the soft wall at $r=1/\sqrt{\mathsf{s}}$. The wall effectively prevents strings from penetrating deep into the bulk.
@@ III Isosceles Triangle Geometry
which are valid for $20\leq\beta\leq(\pi-\theta)/$ 2.
@@ III.1.2 Configuration II
Now we would like to examine the $-\ell$ behavior of $E^{\,\text{\tiny II}}$. As before, the basic idea is to let one or both of the $\lambda$ parameters approach $1$. There are two possible cases to consider: $\lambda_{1}\rightarrow 1$ at fixed $\lambda_{3}<1$, where only strings 1 and 2 (ending on $Q_{1}$ and $Q_{2}$) become infinitely long, and $\lambda_{1},\,\lambda_{3}\rightarrow 1$ , where all the strings become infinitely long.
@@
The key new feature here is that the constant term depends explicitly on the apex angle $\theta$. This has a simple explanation. In the limit as $\lambda_{1}\rightarrow 1$, the two strings become infinitely long, while the third remains finite in size. Its length, as well as the energy, depends on the apex angle that directly affects the constant term but not the leading one.
@@ III.1.3 Configuration III
Finally, consider configuration $III$. A straightforward way to proceed is as follows. Since $\alpha_{3}$ changes its sign from negative to positive, the expressions for the geometrical constraint, leg length, and energy can be obtained from those of subsection 2 by respectively replacing ${\cal L}^{-}$ and ${\cal E}^{-}$ with ${\cal L}^{+}$ and ${\cal E}^{+}$. The force balance equations, however, remain unchanged. Thus, we can rewrite (3.8) and (3.9) as
@@ III.2.1 $\theta<\frac{2}{3}\pi$
The two main conclusions follow from this Table. First, as seen from the fourth column, the $Y$-point generally does not coincide with the Fermat point $F$ of the triangle. [^12] As shown in subsection A, $Y$ approaches $F$ in the limit $\ell\rightarrow\infty$. The only exception is the case $\theta=\frac{\pi}{3}$, where the two points coincide due to symmetry. Second, the last column shows that the coupling $\alpha_{\text{\tiny QQ}}$ depends weakly on $\theta$, and that the ratio $\alpha_{\text{\tiny QQ}}/\alpha_{\text{\tiny Q}\bar{\text{\tiny Q}}}$ is very close to the phenomenologically expected value of $\frac{1}{2}$. Since the discrepancy among the $\alpha_{\text{\tiny QQ}}$ values does not exceed $1.5\%$, the assumption of pairwise quark interaction at small separations appears to be quite reasonable.
@@ IV.2 Some examples
where ${\text{v}}_{\text{\tiny 1}}^{\text{\tiny+}}$ is the solution to Eq.(4.8) (see also Table II).
@@
[…]
@@ IV.3 A remark on diquarks
with $L^{\text{\,\tiny A,\tiny B}}=\frac{1}{\sqrt{\mathsf{s}}}\bigl{(}{\cal L}^{-}(\lambda_{2},v)-{\cal L}^{+}(\alpha_{3},v)\bigr{)}$. The bounds on $v$ are determined by the Equations: $L(v)=\ell$ for $\text{v}_{\ast}$, $\alpha_{1}(v)=0$ for ${\text{v}}_{\text{\tiny 0}}$, and $\lambda_{2}(v)=1$ for ${\text{v}}_{\text{\tiny 1}}$. The second corresponds to a transition between the configurations, while the third to the limit of infinitely long string ending on $Q_{2}$. In contrast to the constraint (4.2), the constraint (4.23) does not prevent us from taking the limit $\lambda_{2}\rightarrow 1$ ($L\rightarrow\infty$) for configuration B.
@@ Appendix C Some details on the quark-antiquark potential
Figure 14: A string configuration in five dimensions. Here, $\ell$ denotes the distance between the two quark sources, and $r_{\text{\tiny 0}}$ is the radial coordinate of the turning point at $x=0$.
@@ D.3 The range $\theta^{1}\leq\theta\leq\theta^{3}$
whose zeros correspond to the solutions of $\alpha_{3}(v)=0$ if $\beta(v)$ and $\alpha_{1}(v)$ are given by the first two equations in (3.33). As before, we are interested in the zeros within the interval $0<v<{\text{v}}_{\text{\tiny 1}}$. A numerical analysis of $f^{3}$ is straightforward and shows two distinct cases: the function has no zeros for $\theta^{1}\leq\theta<0.350\pi$, and has two zeros for $0.350\pi\leq\theta\leq\theta^{3}$. The former case aligns with the earlier discussion of $\theta=\frac{\pi}{3}$, while the latter implies transitions between configurations II and III. For illustration, we plot $f^{3}$ as a function of $v$ in Figure 17. The potential is described in terms of these
@@
configurations as follows: